# ConfusionPrompt: Practical Private Inference for Online Large Language Models

### v4

@@ Abstract
State-of-the-art large language models (LLMs) are typically deployed as online services, requiring users to transmit detailed prompts to cloud servers. This raises significant privacy concerns. In response, we introduce ConfusionPrompt, a novel framework for private LLM inference that protects user privacy by: (i) decomposing the original prompt into smaller sub-prompts, and (ii) generating pseudo-prompts alongside the genuine sub-prompts, which are then sent to the LLM. The server responses are later recomposed by the user to reconstruct the final output. This approach offers key advantages over previous LLM privacy protection methods: (i) it integrates seamlessly with existing black-box LLMs, and (ii) it delivers a significantly improved privacy-utility trade-off compared to existing text perturbation methods. We also develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements for a privacy-preserving group of prompts and provide a complexity analysis to justify the role of prompt decomposition. Our empirical evaluation shows that ConfusionPrompt achieves significantly higher utility than local inference methods using open-source models and perturbation-based techniques, while also reducing memory consumption compared to open-source LLMs.
keywords:
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks [1, 2, 3], leading to their growing integration into real-world applications such as medical consultations [4] and financial services [5]. State-of-the-art LLMs, such as OpenAI’s ChatGPT [6], are predominantly offered as online services. This is mainly due to intellectual property considerations, as model owners are reluctant to disclose their proprietary parameters to users [7]. However, direct usage of such LLM online services could raise significant privacy concerns, as users often transmit prompts containing sensitive information that should ideally remain confidential from the server.
@@
Existing solutions to privacy-preserving model inference, such as encryption [8, 9] and perturbation [10] techniques, are often impractical for private LLM inference. In encryption-based methods, the server leverages homomorphic encryption (HE) [11] and Secure Multiparty Computation (MPC) [12] to conduct private inference on the users’ encrypted query. The practicality of this method is limited by its high computation overheads, especially on LLMs (e.g., GPT-3 has 175B parameters [1]). On the other hand, the basic idea of perturbation-based methods is to inject a specific level of noise into the user’s input before releasing it to the server, which is challenging to strike a satisfied privacy-utility balance [13].
@@
Another limitation of the aforementioned solutions is their dependence on the service provider to modify their infrastructure or disclose specific model parameters, which can be prohibitively costly. Encryption-based techniques require significant investment from platform in hardware and algorithmic acceleration to support efficient computation. Perturbation-based methods often rely on local differential privacy (LDP) [14] to ensure privacy with formal guarantees. For improved LDP performance, recent studies have proposed deploying certain modules on the user side, necessitating the sharing of specific model parameters by the server [15, 16, 17]. These requirements impose additional burdens and proprietary concerns on the service provider, hindering the deployment of these solutions.
@@
To alleviate the above concerns, this paper proposes ConfusionPrompt, a novel private LLM inference framework that can be seamlessly integrated with existing online black-box LLMs. The core idea of ConfusionPrompt is to construct a set of prompts, containing real and fake prompts, designed to confuse the server, making the curious server fail to infer the real user prompt. However, a significant challenge arises when a prompt contains multiple sensitive attributes. In such cases, the user needs to consider all combinations of true and fake attributes to effectively confuse an attacker with prior knowledge of certain attributes, rendering the query complexity growing exponentially with the number of attributes. To address this issue, we propose a decomposition approach in which the query is divided into sub-questions. This ensures that the private information is distributed across different sub-queries, thereby reducing the complexity to approximately linear in the nunber of attributes (see Figure 2).
@@
Building on the aforementioned concepts, our proposed ConfusionPrompt consists of four critical steps: (1) The user decomposes the original prompt into several sub-prompts; (2) For each genuine sub-prompt, the user generates a set of pseudo prompts that obfuscate sensitive attributes in the genuine sub-prompt, continuing this process until the desired level of privacy is achieved. (3) The user sends a group of both genuine and pseudo-prompts to the online service, retrieves the responses, and isolates the sub-responses corresponding to the genuine sub-prompts. (4) Finally, user recomposes the sub-responses to obtain the final result. This framework empowers the user with independent deployment, eliminating the need for additional privacy-preserving actions from the service provider.
@@
  • 1.
    We are the first to propose a private LLM inference framework using confusion-based strategy. Our ConfusionPrompt framework can be seamlessly implemented by clients with existing online black-box LLMs, such as ChatGPT and Claude, providing an improved privacy-utility trade-off. In addition, Our framework can be integrated with Trusted Execution Environment (TEE) [18] solutions. Even in the event of a TEE compromise, user privacy remains safeguarded through the confusion strategy.
  • 2.
    To ensure user privacy, we define a privacy model to formulate the requirements of a group of prompts, including both real and pseudo prompts. Furthermore, we introduce a local decomposition module to reduce the complexity of the prompt group under the same privacy parameters.
  • 3.
    Experiments show that ConfusionPrompt can achieve a consistently and significantly better utility than the local inference method using open-source models and LDP-based methods. Moreover, ConfusionPrompt significantly reduces memory overhead compared to open-source LLMs.
@@ 2 Related Works
In this section, we review three types of methods for privacy-preserving LLM inference: anonymization-based method, encryption-based method, and perturbation-based method.
@@
Anonymization-based method. Traditional anonymization techniques rely on Named Entity Recognition (NER) to redact Personally Identifiable Information (PII) such as individuals’ names, social security number, and email addresses [19, 20]. Recent research on LLM inference involves sanitizing sensitive items in the input and subsequently de-anonymize the LLM’s returned responses [21, 22]. The downside of this method is that anonymization fall short in concealing other sensitive information, including verbs and non-named entities [23].
@@
Encryption-based method. Cryptonets [24] proposed the first neural network inference on encrypted data using homomorphic encryption (HE). They approximated the non-linear function such as Sigmoid and MaxPooling by polynomials. Iron [25] designed specialized and efficient protocols for two types of computationally heavy operations in Transformer-based inference: (i) matrix multiplications, and (ii) complex functions including Softmax, GELU activations, and LayerNorm. To achieve further speedup, [8] transformed the high-overhead functions into cryptography-friendly approximations, and finetuned the model to maintain accuracy.
@@
Perturbation-based method. Perturbation-based methods provide privacy guarantee by inject calibrated noise into the input. Existing studies utilizing this method predominantly focus on privacy protection in fine-tuning [26, 27] and prompt-tuning phases [28, 29], while few studies investigate the privacy-preserving inference paradigm. A major challenge in perturbation-based private inference is to balance the utility and privacy tradeoff. Recent studies have proposed Text2Text [30] and paraphraser-based approaches [31, 32] to privatize text with LDP. Split-N-Denoise (SnD) [16] deployed the token embedding layer at the client side, and introduced a user-side denoising model to correct the purturbed embedding output for downstream tasks.
@@ 3 ConfusionPrompt: Design and Formulation
In this section, we first illustrate the overview of our proposed ConfusionPrompt, which consists of three critical components: decomposer, generator, and recomposer. Then, we develop a privacy model to formulate the requirements for achieving privacy-preserving inference.
@@ 3.1 Overall Design
Denote $G_{s}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ as the LLM deployed as cloud service such as ChatGPT [6, 7], where $\mathcal{V}^{*}$ is the vocabulary space. Instead of sending raw user prompt to the cloud, our ConfusionPrompt introduces a new privacy-preserving inference paradigm, which makes it significantly hard for any third party to infer the real original user prompt even if it happens to access the information sent from the user. Basically, ConfusionPrompt involves the following six steps:
@@
Step 3: evaluation of privacy level. To ensure that the genuine prompt is safely hidden in the prompt group, we also need to design some criteria (e.g., semantic irrelevance) to evaluate the usability of generated pseudo prompts. Based on these criteria, we iteratively sample and filter the pseudo prompts until the prompt group meets the privacy requirement, which will be introduced in details in Section 3.2.
Step 4: communication with cloud server. Here, the user sends the prompt group to the server for the usage of LLM on the cloud, while the server subsequently sends back the response group to the user. During this process, we expect that it will be hard for the server to infer the original user prompt according to the prompt group, therefore safeguarding user’s privacy.
Step 5: retrieval of interested response. Since the user itself is aware of which sub-prompts are genuine, it can seamlessly retrieve the corresponding sub-responses and discard the responses of pseudo prompts.
@@ 3.2.1 Rationale of Privacy Model
Here, we define a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements that the group of pseudo prompts should satisfy for privacy protection, which offers guidance for the evaluation of privacy level and training of models (i.e., decomposer, generator, and recomposer). To explain the rationale of our privacy model, we follow [33] to quantify the privacy risk of the queries exposed to the server.
@@
Equation 2 suggests three possible methods to minimize the privacy risk. First, reducing $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$ for any $\boldsymbol{p}^{\prime}\neq\boldsymbol{p}^{0}$ leads to a decrease in $\mathrm{Risk}(\boldsymbol{P})$. Therefore, generating deceptive prompts that are semantically distinct from the genuine prompt can effectively lower the privacy risk. Second, $\mathrm{Risk}(\boldsymbol{P})$ is reduced when $\pi(\boldsymbol{p}^{\prime}|\boldsymbol{P})$ is high for prompts $\boldsymbol{p}^{\prime}$ that do not resemble $\boldsymbol{p}^{0}$. This implies that fabricated prompts should be crafted to appear as realistic as possible. Finally, $\mathrm{Risk}(\boldsymbol{P})$ decreases if we increase the number of prompts $\boldsymbol{p}^{\prime}$ with low $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$. This can be achieved by increasing the number of fake prompts with different semantic meaning.
Based on the observations, we develop the following criteria for the fabricated prompts: (i) The fabricated attributes in the pseudo prompt should be semantically irrelevant to the original attributes in genuine prompt, so that the genuine user information can be effectively obfuscated [34]. (ii) The genuine prompt should be obfuscated by sufficient number of pseudo prompts, and the attacker can not easily identify the real prompt from the combination pattern even with background knowledge. (iii) The pseudo prompt should appear genuine rather than obviously fabricated. Thus it’s crucial to maintain the fluency and reasonability of the pseudo prompts.
@@ 3.2.2 Construction of Privacy Model
Given two attributes $u_{1}$, $u_{2}\in\mathcal{U}$, an attribute similarity function measures the semantic similarity between $u_{1}$ and $u_{2}$: $\rm{Sim}(u_{1},u_{2}):\mathcal{U}\times\mathcal{U}\rightarrow\mathbb{R}$.
For a pair of genuine and pseudo prompts, we define the correspondent attributes as followed:
@@
Given two prompts $\boldsymbol{p}_{1}$, $\boldsymbol{p}_{2}\in\mathcal{P}$, the correspondent attribute of $u_{i}\in U(\boldsymbol{p}_{1})$ in $\boldsymbol{p}_{2}$, is defined as the attribute at the same syntactic position of $\boldsymbol{p}_{2}$, denoted by $\rm{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$.
@@
$\mathrm{Sim}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=\max_{u_{i}\in U(\boldsymbol{p}_{1})}\rm{Sim}\left(u_{i},\rm{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\right)$ (3)
We proceed to the definition of significance of an attribute, which is tied to the likelihood of a curious server being able to identify the genuine attribute from a group of genuine and pseudo-prompts.
@@
$H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})=\left\{\begin{array}[]{ll}1&\rm{Corr}\left(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j}\right)=u\\ 0&\rm{otherwise}\end{array}.\right.$ (5)
In our setting, the group of prompts consists of one genuine prompt along with a collection of pseudo prompts. The significance of a true attribute can be considered as the proportion of its occurrence within the group of prompts. In the following, we provide the definition of significance for multiple attributes.
@@
$p_{ih}$ is the $h$-th decomposed sub-prompt in the $i$-th prompt. In case that the attacker have prior knowledge of some private attributes, we introduce a notation $\mathcal{V}$ as a subset of private attributes in $p_{0h}$. Then we consider the proportion of a given attribute conditioning on any possible set of private attributes to bound the possibility of correctly identify the target attribute with some background knowledge. Now, we provide the definition of genuineness for the third criterion:
@@
Denote $D:\mathcal{P}\rightarrow\mathbb{R}$ as a function to discriminate between fabricated and genuine prompts. A larger value returned by $D$ indicates a higher likelihood of a prompt being genuine. The genuineness of a prompt $p\in\mathcal{P}$ can be defined as $\rm{Genu}(p)=D(p)$.
@@
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are pseudo prompts. If $\boldsymbol{P}$ satisfies the following requirements, then it is deemed that it can ensure $(\lambda,\mu,\rho)$-privacy of user prompt $\boldsymbol{p}_{0}$ with respect to discriminator $D$:
  • 1.
    Each pseudo prompt should be semantically irrelavant to user prompt $\boldsymbol{p}_{0}$, i.e., $\forall i\in[1,n]$, $\rm{Sim}(\boldsymbol{p}_{0},\boldsymbol{p}_{i})\leq\lambda$.
  • 2.
    The users’ sensitive attributes should be obfuscated by sufficient pseudo prompts, i.e., $\rm{Sig}(U(\boldsymbol{p}_{0}),\boldsymbol{P})\leq\mu$.
  • 3.
    Each pseudo prompt should not be classified as fabricated prompt by discriminator $D$, i.e., $\forall i\in[1,n]$, $\rm{Genu}(p_{i})=D(p_{i})\geq\rho$.
In Appendix B, we provide the relationship between our privacy model and the upper bound on the success rate of inference attacks, which states that an curious server can do no better than random guessing if the combination of attributes are uniformly distributed.

### v5

@@ Abstract
State-of-the-art large language models (LLM) are often deployed as online services, posing privacy risks for user prompts. In response, we introduce ConfusionPrompt, a novel framework for private LLM inference that protects user privacy by: (i) decomposing the original prompt into smaller sub-prompts, and (ii) generating pseudo-prompts alongside the genuine sub-prompts. ConfusionPrompt seamlessly integrates with existing black-box LLMs and achieves a superior privacy-utility trade-off compared to existing text perturbation methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements for a privacy-preserving group of prompts and provide a complexity analysis to justify the role of prompt decomposition. Our empirical evaluation shows that ConfusionPrompt outperforms local inference with open-source models by over 11% and surpasses perturbation-based techniques by over 42%, while reducing memory consumption by at least 62% compared to open-source LLMs.
{IEEEImpStatement}
The rapid adoption of LLMs via cloud-based services has heightened privacy concerns, as users must transmit sensitive prompts to external servers. Existing approaches for privacy-preserving LLM inference, such as encryption-based techniques and perturbation-based strategies, are often impractical as they either significantly degrade model performance or lack compatibility with black-box LLMs. Our proposed framework, ConfusionPrompt, introduces a novel, client-side framework that enables private inference without requiring any modifications to existing black-box LLM services. ConfusionPrompt consistently outperforms local inference with open-source LLMs and perturbation-based privacy methods in terms of utility, while also minimizing memory and computational overhead. This framework advances toward safer, privacy-conscious deployment of LLMs, an increasingly pivotal concern in real-world AI applications.
{IEEEkeywords}
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks [6, 55, 28, 7, 68], driving their adoption in real-world applications such as medical consultations and financial services [67, 66, 59]. State-of-the-art LLMs, such as OpenAI’s ChatGPT [45], are predominantly offered as online services, primarily to safeguard proprietary model parameters and intellectual property [44]. However, this paradigm introduces significant privacy risks, as users often transmit prompts containing sensitive information that should ideally remain confidential from the server.
@@
Existing solutions to privacy-preserving model inference, such as encryption [37, 9] and perturbation [15] techniques, are often impractical for private LLM inference. In encryption-based methods, the server leverages homomorphic encryption (HE) [1] and Secure Multiparty Computation (SMPC) [14] to conduct private inference on the users’ encrypted query. The practicality of this method is limited by its high computation overheads, especially on LLMs (e.g., GPT-3 has 175B parameters [6]). On the other hand, the basic idea of perturbation-based methods is to inject a specific level of noise into the user’s input before releasing it to the server, which is challenging to strike a satisfied privacy-utility balance [40].
@@
Another limitation of the aforementioned solutions is their dependence on the service provider to modify their infrastructure or disclose specific model parameters, which can be prohibitively costly. Encryption-based techniques require significant investment from platform in hardware and algorithmic acceleration to support efficient computation. Perturbation-based methods often rely on local differential privacy (LDP) [18] to ensure privacy with formal guarantees. For improved LDP performance, recent studies have proposed deploying certain modules on the user side, necessitating the sharing of specific model parameters by the server [50, 41, 42, 34]. These requirements impose additional burdens and proprietary concerns on the service provider, hindering the deployment of these solutions. It remains an open challenge to develop a method that seamlessly integrates with existing black-box LLMs while maintaining strong utility.
@@
To alleviate the above concerns, this paper proposes ConfusionPrompt, a novel private LLM inference framework that can be seamlessly integrated with existing online black-box LLMs. The core idea of ConfusionPrompt is to construct a set of prompts, containing real and fake prompts, designed to confuse the server, preventing the server from accurately identifying the genuine user prompt. However, a significant challenge arises when a prompt contains multiple sensitive attributes. In such cases, the user needs to consider all combinations of true and fake attributes to effectively confuse an attacker with prior knowledge of certain attributes, rendering the query complexity growing exponentially with the number of attributes. To address this issue, we propose a decomposition approach in which the query is divided into sub-questions. This ensures that the private information is distributed across different sub-queries, thereby reducing the complexity to approximately linear in the number of attributes (see Figure 2).
@@
Building on the aforementioned concepts, our proposed ConfusionPrompt consists of four critical steps: (1) The user decomposes the original prompt into several sub-prompts. (2) For each genuine sub-prompt, the user generates a set of pseudo prompts that obfuscate sensitive attributes in the genuine sub-prompt, repeating this process until the desired level of privacy is achieved. (3) The user sends a group of both genuine and pseudo-prompts to the online service, retrieves the responses, and isolates the sub-responses corresponding to the genuine sub-prompts. (4) Finally, user recomposes the sub-responses to obtain the final result. This framework allows for independent user-side deployment, eliminating the need for privacy-preserving modifications by the service provider.
@@
(1) We are the first to propose a private LLM inference framework using confusion-based strategy. Our ConfusionPrompt framework can be seamlessly implemented by clients with existing online black-box LLMs, such as ChatGPT and Claude, providing an improved privacy-utility trade-off. In addition, our framework can be integrated with Trusted Execution Environment (TEE) [56] solutions. Even in the event of a TEE compromise, user privacy remains safeguarded through the confusion strategy.
(2) To ensure user privacy, we define a privacy model to formulate the requirements for a group of prompts, including both real and pseudo-prompts. Furthermore, we introduce a local decomposition module to reduce the complexity of the prompt group under the same privacy parameters.
(3) Experiments show that ConfusionPrompt can achieve a consistently and significantly better utility than LDP-based methods. Moreover, ConfusionPrompt surpasses the performance of local inference method using open-source models with lower memory requirement.
@@ 2 Related Works
In this section, we review three types of methods for privacy-preserving LLM inference: sanitization-based method, encryption-based method, and perturbation-based method.
@@
Sanitization-based method. Traditional sanitization techniques rely on Named Entity Recognition (NER) to redact Personally Identifiable Information (PII) such as individuals’ names, social security number, and email addresses [19, 32, 48]. The downside of this method is that NER falls short in concealing other sensitive information, including verbs and non-named entities [3]. Recent research on LLM inference involves sanitizing sensitive items in the input and subsequently de-anonymize the LLM’s returned responses [29, 10, 49]. However, such input perturbations can introduce semantic biases, potentially leading to responses that are semantically irrelevant. Casper utilizes a combination of rule-based and machine learning (ML)-based techniques to filter sensitive PIIs and alert users of the sensitive topics [12]. While this method effectively detects sensitive information, it does not address how to process such information without adversely affecting the output of LLMs. In general, sanitization approaches identify and remove sensitive information, but this can degrade output utility when such information is crucial for LLM inference.
@@
Encryption-based method. Cryptonets [23] proposed the first neural network inference on encrypted data using homomorphic encryption (HE). They approximated the non-linear function such as Sigmoid and MaxPooling by polynomials. Iron [24] designed specialized and efficient protocols for two types of computationally heavy operations in Transformer-based inference: (i) matrix multiplications, and (ii) complex functions including Softmax, GELU activations, and LayerNorm. To achieve further speedup, [37] transformed the high-overhead functions into cryptography-friendly approximations, and finetuned the model to maintain accuracy. SecFormer designs a combination of SMPC protocols for accurate and efficient computation of complex nonlinear functions, such as GeLU and LayerNorm [39]. BOLT reduces the payload on both linear and non-linear operations through cryptographic optimizations, and eliminates insignificant words using ML techniques for enhanced efficiency [47]. NEXUS introduces the first non-interactive protocol, allowing the client to complete the entire inference process with the server in a single round of communication [72]. Despite their privacy advantages, encryption-based methods impose significant computational overhead and require substantial investment in hardware and algorithmic acceleration, making integration with existing black-box LLMs challenging.
@@
Perturbation-based method. Perturbation-based methods provide privacy guarantee by injecting calibrated noise into the input. Existing studies utilizing this method predominantly focus on privacy protection in fine-tuning [31, 71, 35] and prompt-tuning phases [17, 36], while few studies investigate the privacy-preserving inference paradigm. A major challenge in perturbation-based private inference is to balance the utility and privacy trade-off. Recent studies have proposed Text2Text [21] and paraphraser-based approaches [64, 43] to privatize text with LDP. Split-N-Denoise (SnD) [41] deployed the token embedding layer at the client side, and introduced a user-side denoising model to correct the purturbed embedding output for downstream tasks. More recently, DYNTEXT proposes a semantic-aware dynamic text sanitization mechanism for privacy-preserving LLM inference, which adaptively perturbs sensitive tokens according to their semantic density to improve the privacy-utility trade-off [73]. InferDPT further studies privacy-preserving inference for closed-box LLMs by combining a perturbation module with a local extraction module, such that a perturbed prompt is uploaded to the remote LLM and a lightweight local model reconstructs a higher-quality output from the perturbed generation [61]. In a related but distinct setting, FedCoT proposes a federated chain-of-thought distillation framework that transfers knowledge from a server-side LLM to a client-side small language model using perturbed prompts and rationales to protect user privacy [20]. However, existing perturbation-based methods either significantly degrade utility or are incompatible with black-box LLMs due to proprietary and efficiency concerns.
In summary, existing approaches either significantly degrade model performance or lack compatibility with black-box LLMs. To address this, we propose a confusion-based strategy that uploads a mix of genuine and synthetic prompts while preserving privacy. Table 1 summarizes the coarse-grained comparision between ConfusionPrompt and existing approaches.
Table 1: Coarse-grained comparison between privacy-preserving LLM inference frameworks.
Method Plug-and-Play Impact on Utility
HaS [10] Yes Medium
Casper [12] Yes Medium
Iron [24] No Low
SecFormer [39] No Low
BOLT [47] No Low
NEXUS [72] No Low
Text2Text [21] Yes High
Paraphraser [64, 43] Yes High
DYNTEXT [73] Yes Medium
InferDPT [61] Yes Medium
SnD [41] No Medium
ConfusionPrompt Yes Low
  ConfusionPrompt offers a plug-and-play solution with little impact on utility. It achieves superior privacy-utility trade-off compared to sanitization and perturbation methods, which often introduce semantic bias into the original prompt.
@@ 3 ConfusionPrompt: Design and Formulation
In this section, we first illustrate the overview of our proposed ConfusionPrompt, which consists of three critical components: decomposer, generator, and recomposer. Then, we develop a privacy model to formulate the requirements for privacy-preserving inference.
@@ 3.1 Overall Design
Denote $G_{s}:\mathcal{V}^{*}\rightarrow\mathcal{V}^{*}$ as the LLM deployed as cloud service such as ChatGPT [45, 44, 8], where $\mathcal{V}^{*}$ is the vocabulary space. Instead of transmitting raw user prompt to the cloud, ConfusionPrompt introduces a privacy-preserving inference paradigm that significantly hinders third parties from inferring the original user prompt, even if they access the transmitted data. Our framework consists of six key steps:
@@
Step 3: evaluation of privacy level. To ensure that the genuine prompt is safely hidden in the prompt group, we also need to design some criteria (e.g., semantic irrelevance) to evaluate the usability of generated pseudo-prompts. Based on these criteria, we iteratively sample and filter the pseudo-prompts until the prompt group meets the privacy requirement, which will be introduced in details in Section 3.2.
Step 4: communication with cloud server. The user transmits the prompt group to the server for LLM processing, and the server returns the corresponding response group. Throughout this process, the prompt group is designed to obscure the original user prompt, making it difficult for the server to infer the sensitives information, thereby safeguarding user’s privacy.
Step 5: retrieval of interested response. Since the user knows which sub-prompts are genuine, they can effortlessly extract the corresponding sub-responses while discarding those from pseudo-prompts.
@@ 3.2.1 Rationale of Privacy Model
In this section, we define a $(\lambda,\mu,\rho)$-privacy model to formulate the requirements that the group of pseudo-prompts should satisfy for privacy protection. This model provides guidance for evaluating privacy levels and training components (i.e., decomposer, generator, and recomposer). To explain the rationale of our privacy model, we follow [46] to quantify the privacy risk of queries exposed to the server.
@@
Equation 2 suggests three possible methods to minimize the privacy risk. First, reducing $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$ for any $\boldsymbol{p}^{\prime}\neq\boldsymbol{p}^{0}$ leads to a decrease in $\mathrm{Risk}(\boldsymbol{P})$. Therefore, generating deceptive prompts that are semantically distinct from the genuine prompt can effectively lower the privacy risk. Second, $\mathrm{Risk}(\boldsymbol{P})$ is reduced when $\pi(\boldsymbol{p}^{\prime}|\boldsymbol{P})$ is high for prompts $\boldsymbol{p}^{\prime}$ that do not resemble $\boldsymbol{p}^{0}$. This implies that fabricated prompts should be crafted to appear as realistic as possible. Finally, $\mathrm{Risk}(\boldsymbol{P})$ decreases if we increase the number of prompts $\boldsymbol{p}^{\prime}$ with low $\mathrm{Sim}(\boldsymbol{p}^{\prime},\boldsymbol{p}^{0})$. This can be achieved by increasing the number of fake prompts with different semantic meanings.
Based on the observations, we develop the following criteria for the fabricated prompts: (i) The fabricated attributes in the pseudo prompt should be semantically irrelevant to the original attributes in genuine prompt, so that the genuine user information can be effectively obfuscated [69]. (ii) The genuine prompt should be obfuscated by sufficient number of pseudo-prompts, and the attacker can not easily identify the real prompt from the combination pattern even with background knowledge. (iii) The pseudo-prompt should appear genuine rather than obviously fabricated. Thus it is crucial to maintain the fluency and reasonability of the pseudo-prompts.
@@ 3.2.2 Construction of Privacy Model
Given two attributes $u_{1}$, $u_{2}\in\mathcal{U}$, an attribute similarity function measures the semantic similarity between $u_{1}$ and $u_{2}$: $\text{Sim}(u_{1},u_{2}):\mathcal{U}\times\mathcal{U}\rightarrow\mathbb{R}$.
For a pair of genuine and pseudo-prompts, we define the correspondent attributes as follows:
@@
Given two prompts $\boldsymbol{p}_{1}$, $\boldsymbol{p}_{2}\in\mathcal{P}$, the correspondent attribute of $u_{i}\in U(\boldsymbol{p}_{1})$ in $\boldsymbol{p}_{2}$, is defined as the attribute at the same syntactic position of $\boldsymbol{p}_{2}$, denoted by $\text{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$.
@@
$\mathrm{Sim}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=\max_{u_{i}\in U(\boldsymbol{p}_{1})}\text{Sim}\left(u_{i},\text{Corr}(u_{i},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\right)$ (3)
Next, we define the significance of an attribute, which is tied to the likelihood of a curious server being able to identify the genuine attribute from a group of genuine and pseudo-prompts.
@@
$H(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j})=\left\{\begin{array}[]{ll}1&\text{Corr}\left(u,\boldsymbol{p}_{i},\boldsymbol{p}_{j}\right)=u\\ 0&\text{otherwise}\end{array}.\right$. (5)
In our setting, the group of prompts consists of one genuine prompt along with a collection of pseudo-prompts. The significance of a true attribute can be considered as the proportion of its occurrence within the group of prompts. In the following, we provide the definition of significance for multiple attributes.
@@
$p_{ih}$ is the $h$-th decomposed sub-prompt in the $i$-th prompt. In case that the attacker have prior knowledge of some private attributes, we introduce a notation $\mathcal{V}$ as a subset of private attributes in $p_{0h}$. Then we consider the proportion of a given attribute conditioned on any possible set of private attributes to bound the possibility of correctly identifying the target attribute with some background knowledge. Now, we provide the definition of genuineness for the third criterion:
@@
Denote $D:\mathcal{P}\rightarrow\mathbb{R}$ as a function to discriminate between fabricated and genuine prompts. A larger value returned by $D$ indicates a higher likelihood of a prompt being genuine. The genuineness of a prompt $p\in\mathcal{P}$ can be defined as $\text{Genu}(p)=D(p)$.
@@
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are pseudo-prompts. If $\boldsymbol{P}$ satisfies the following requirements, then it is deemed that it can ensure $(\lambda,\mu,\rho)$-privacy of user prompt $\boldsymbol{p}_{0}$ with respect to discriminator $D$:
  • Each pseudo prompt should be semantically irrelavant to user prompt $\boldsymbol{p}_{0}$, i.e., $\forall i\in[1,n]$, $\text{Sim}(\boldsymbol{p}_{0},\boldsymbol{p}_{i})\leq\lambda$.
  • The users’ sensitive attributes should be obfuscated by sufficient pseudo-prompts, i.e., $\text{Sig}(U(\boldsymbol{p}_{0}),\boldsymbol{P})\leq\mu$.
  • Each pseudo-prompt should not be classified as fabricated prompt by discriminator $D$, i.e., $\forall i\in[1,n]$, $\text{Genu}(p_{i})=D(p_{i})\geq\rho$.
In .2, we provide the relationship between our privacy model and the upper bound on inference attacks.
3.2.3 Implication and Selection of Privacy Parameters
In this section, we provide a explanation for the three parameters in our privacy model:
  • Similarity parameter $\lambda$ controls the similarity between private attributes in genuine and pseudo-prompts. A lower $\lambda$ results in more dissimilar pseudo-prompts, thereby enhancing privacy protection. In our empirical analysis, we measure cosine similarity in the embedding space of private attributes. As shown in Figure 4, reducing $\lambda$ decreases the success rate of attribute inference attacks, with attack accuracy remaining sufficiently low for $\lambda<0.8$ under appropriate choices of the other two parameters (i.e., $\mu>10$ and $\rho=4$).
  • Significance parameter $\mu$ provides the bound on attack success rate of correctly identifying the target attribute. A lower $\mu$ necessitates generating a larger number of pseudo-prompts to obscure sensitive information, while providing higher level of privacy protection. The choice of $\mu$ depends on the user’s privacy requirements and the attribute space conditioned on the non-private information.
  • Genuineness parameter $\rho$ quantifies the degree of linguistic realism exhibited by individual synthetic prompts. In empirical analysis, we define the genuineness of a prompt on a four-point ordinal scale: 1 (incomprehensible), 2 (low quality), 3 (moderate fluency), and 4 (perfect fluency). We recommend the user to configure $\rho=4$, i.e., requiring each pseudo-prompt to satisfy perfect fluency, so that the attacker can not easily detect fabricated prompts.